# Sample Observed Effects: Enumeration, Randomization and Generalization

### v5

@@ Abstract
The widely used ’Counterfactual’ definition of Causal Effects was derived for unbiasedness and accuracy - and not generalizability. We propose a Combinatorial definition for the External Validity (EV) of intervention effects. We first define the concept of an effect observation ’background’. We then formulate conditions for effect generalization based on their sets of (observed and unobserved) backgrounds. This reveals two limits for effect generalization: (1) when effects are observed under all their enumerable backgrounds, or, (2) when backgrounds have become sufficiently randomized. We use the resulting combinatorial framework to re-examine several issues in the original counterfactual formulation: out-of-sample validity, concurrent estimation of multiple effects, bias-variance tradeoffs, statistical power, and connections to current predictive and explaining techniques.
Methodologically, the definitions also allow us to replace the parametric estimation problems that followed the counterfactual definition by combinatorial enumeration and randomization problems in non-experimental samples. We use this non-parametric framework to demonstrate (External Validity, Unconfoundness and Precision) tradeoffs in the performance of popular supervised, explaining, and causal-effect estimators. We also illustrate how the approach allows for the use of supervised and explaining methods in non-i.i.d. samples. The COVID19 pandemic highlighted the need for learning solutions to provide predictions in severally incomplete samples. We demonstrate applications in this pressing problem.
@@ 1 Introduction
Non-parametric effect estimators (such as permutation tests  [Hoeffding, 1948] and feature importance attribution  [Burkart and Huber, 2021]) can often only be used under strict assumptions of no unobserved factors. Consider, for example, a sample containing units that have received a treatment (’treatment’ units) and others that did not (’control’). To estimate effects, a simple strategy is to permute treatment assignment among all treatment-control pairs that are similar enough across relevant sample variables  [Morgan and Winship, 2007]. This strategy relies, however, on the assumption that paired units are exchangeable besides their treatment status. The assumption is easily violated if units differ in any unobserved way (even in high-dimensional samples it is hard for researchers to make the claim that there are no overlooked, relevant distinctions). The main challenge is that of unobserved confounding. Confounding refers to sampling conditions where uncontrolled variation of out-of-sample factors bias in-sample estimates. Under these conditions, sample observation of effects, for example, reflect not only the true effect of sample factors, but also of uncontrolled and unobserved factors. In Experimental Design, randomization plays a crucial role in alleviating the effects of unobserved confounding  [Zhang and Zhao, 2023, Kempthorne and Doerfler, 1969]. Randomized Experiments provide a solid foundation, and are still the golden rule, for making causal inferences about the effects of treatments in the Sciences. Non-experimental samples can be seen as experimental samples where randomization did not occur perfectly. Despite these known shortcomings, permutation and Shapley-value based tests  [Burkart and Huber, 2021, Ribeiro, 2022a], continue to be used in practice in non-experimental and partially observed data, with little clarity over the resulting effect estimates’ generalizability and biases. We provide a combinatorial definition for effect generalizability and demonstrate how to select subsamples in non-experimental data that lead to effect estimates with high generalizability and low confoundness. We demonstrate uses of these concepts when estimating causal effects non-parametrically or when explaining the performance of supervised predictors. The latter is currently an intensely studied subject in Machine Learning and Artificial Intelligence  [Lundberg and Lee, 2017, Burkart and Huber, 2021].
@@
A sample’s Data Generating Process (DGP) refers to the underlying mechanism or model that generated the sample. Partition the set $Z$ of DGP variables into two subsets, $Z=X_{m}\cup U_{q}$, the first subset with its $m$ observed variables $X_{m}=\{-1,+1\}^{m}$, and the second with $q$ unobserved $U_{q}=\{-1,+1\}^{q}$. Imagine also variables in $X$ and $U$ are constrained by precedence relationships, as described by two partially ordered set (posets) $G(Z)=\{G(X),G(U)\}$. The distinct ways in which $\{G(X),G(U)\}$ can be combined is described by the distinct linear extensions among these posets. For poset chains $\{G(X),G(U)\}=\{\;\mathbf{a}<\mathbf{b},\;\;u_{1}<u_{2}<u_{3}\;\}$, for example, the extensions are
@@
The different extensions can be seen as different confounding conditions, and orders, under which sample $X$ can be observed. We imagine real-world samples can be collected under any such conditions, and we would like to formulate statistical properties for effects across partially observed samples, $q>0$. The likelihood of a $U{-}X$ extension follow, in turn, relationships between the DGP probability distributions for $X$ and $U$,
@@ 1.0.1 Alternative Orders and Confounding: A Preliminary illustration
For illustration, consider the observed effects of sample factors in a sample with a true cause $a$, a confounder $b$ (with effects on $y$ fully correlated with its root cause $a$), and a spurious variable $c$, $X=\{a,b,c\}$. We will use Latin square diagrams (’squares’) as aid to illustrate several concepts in this article. Fig.1(a) shows a $4$-factor square. Each square is associated with a sample unit $x_{0}$, its reference unit, which stratifies the sample (i.e., other units) across its cells. A square with reference $x_{0}$ correspond to the set of effect observations with singleton factor differences,
$\begin{split}\bigcup_{x\in\mathcal{P}(X)}{{\Delta}}y(x-x_{0})=\{a\},\qquad\forall\,a\in X.\end{split}$
where $\mathcal{P}(X)$ is the powerset (set of sets) of $X$ and ${\Delta}y(x-x_{0})$ is an observation of effect in an outcome-of-interest, $y\in\mathbb{R}$, after an intervention on factors $x-x_{0}$. A complete square contains therefore $\binom{m}{2}$ effect observations, and a sample (or all its squares) is limited to $2^{m}$ observations. We will say that the square’s diagonals enumerate, for example, observations of effect for fixed factors across different (observed) ’backgrounds’. For the cause $a$ and confounder $b$, effect observations are therefore
@@
We see that when the cause $a$ is in the sample, it sustains its effects through arbitrary sample imputations (indicated by the conditional statements), while the confounder also assumes negative values. In a square with $X$ factors, observed under different orders, effect observations for confounders change many times, while causes remain invariant throughout the observed range of background variation.
@@
For any factors $a,b\in Z$, observing their effects under conditions $\Delta y(a\,|\,b)$ and $\Delta y(b\,|\,a)$, or in ordinal relation $a>b$ and $b>a$, is necessary to identify, from effect observations, whether $a$ confounds $b$ (and vice versa). When observed under such conditions, effect variances, $\textrm{Var}[\Delta\hat{y}(b)]$, for causes, confounders and spurious factors become distinct, and have expected values 00, $\nicefrac{{\rho_{ab}}}{{4}}\Delta y(a)$, $\nicefrac{{1}}{{2}}\Delta y(a)$, where $\rho_{ab}$ is the Spearman’s rank correlation coefficient between confounder and its root cause, and $\Delta y(a)$ the root cause’s effect.
According to the principle, under specific ordinal conditions on $U{-}X$ factors, effect observations for confounders have an expected effect variance of $0.25\times\rho_{ab}$ (i.e., the number of times the confounder appears before its root cause in square rows) and causal variables 00 of the true effect $\Delta y(a)$. One such condition is of a completely observed sample, $q=0$. In this case, the principle is associated with the observation of each effect under the cyclic permutation of all others factors (as illustrated by distinct square rows). The principle then states that effect variances, ${\rm I\kern-1.99997ptE}[\,\textrm{Var}[\Delta\hat{y}(a)]\,]$, are then a sufficient statistic to identify confounders under a fixed level of cause-confounder correlation $\rho_{ab}$. Other conditions, $q>0$, are established below.
@@ 1.1 Effect estimates and Unobserved Sample Variation
Before formulating necessary conditions for the generalizability and confoudness of effect observation in generic samples, we formulate our working definition of effects, their generalizability, and a roadmap to the article. According to the Potential Outcomes framework’ definitions for what is a causal effect  [Morgan and Winship, 2007, Rubin, 2005], if $y$ is an outcome of interest and $a$ is a treatment indicator, then the causal effect of $a$ is the difference
@@
where $y_{\leq}$ is the observed outcome $y$ of a sample unit with, exclusively, the set of factors before $a$ in the permutation order $\pi$, and $y_{>}$ of units also containing the factors after $a$. Difference between units in Eq.(3) constitutes observations of effects of the same factor in distinct backgrounds, $\pi\in\Pi_{n}(Z/\{a\})$. Each background fixes a set of factors for their ’treatment’ (before $a$ in $\pi$ or, similarly, sample effects that do not reflect the effect of $a$) and ’control’ (after $a$) for sample units.
Furthermore, $\Pi_{n}(Z/\{a\})$ denotes a subset of the $(m+q)!$ full set of permutations of factors $Z$ observable between sequential changes in $a$ values. What permutation subsets can offer guarantees over the generalizability of sample effect estimates? We formulate the concept of External Validity increasing (EV-increasing) permutations as those that can increase the generalizability of effects in samples. The article’s main argument is structured as follows,
@@
EV is here the inverse of the variance (i.e., the precision) of observed effects under all (e.g., asymptotically) EV-increasing backgrounds, $\Pi_{n}(Z/\{a\})$, in an $U-X$ system for a treatment $a$. This defines effects in a way that is almost opposite to Eq. (1). It calls for effects to be observed under large variation, as opposed to no variation. According to Eq.(3,4), the counterfactual ideal in Eq.(1) is, in fact, the estimate with minimal EV. Underlying these definitions and Principle (1) is the observation that, unlike confounders, causes can be fundamentally defined as ’perfect controls’  [Pearl, 2000]. Similar definitions have been explored in model robustness research  [Peters et al., 2016, Buehlmann, 2020]. The non-parametric definitions here frame the problem, in particular, not as simply the invariance (of effects), but invariance of effects under maximum variance of their ’backgrounds’. Shifting the focus to the later, $\Pi_{n}(Z/\{a\})$, is a key but challenging part of the proposed solution.
Figure 1: (a) $4{\times}4$ Latin-Square (’square’) as sets of effect observation for a sample unit $x_{0}\subseteq X$, (b) $3$ samples (horizontal lines, top) with increasing dimension $m=\{1,2,4\}$ and their ordinal statistics (interval ticks) , as well as alternative samples of confounding unobserved factors $U$ (circles)(top), resulting distribution and percentage of enumerable permutations of effect observations across ordinal variables (bottom and right).
@@ 2 Ordinal Statistics and Sample Confounding Gaps
If $\{X_{a},X_{b},\ldots,X_{m}\}$ are random samples from a particular DGP distribution on $X_{m}$, it will be important to know whether changes in a factor $a$ are expected to follow or precede changes in another factor $b$. The likelihood of such order relations is determined by the DGP, as illustrated in Example. (2). To describe such patterns more formally, we define $X_{(1)},X_{(2)},\ldots,X_{(m)}$ to be an ordered version of sample factors  [Gut, 2009], that is,
@@
Notice that $X_{\min}$ is a random variable. For simplicity, we assume that, in sufficiently large samples and time resolution, ranks do not coincide, leading to strict inequalities (as opposed to $\leq$).
Suppose that $X_{1},X_{2},\ldots,X_{m}$ follow a continuous distribution with PDF $p$ and CDF $F$. We will use $X$, $U$ or $Z$ as alternative labels when considering distributions over observed, unobserved or either factors. Imagine, in particular, a random sample of $U$ factors when their minimum is below an in-sample value $x$. Fig.1(b) give examples where circles are $U$ samples and interval ticks are $X$ samples. The successive lines show $X_{(k)}$ in samples with increasing dimensionality $m$. Circles to the left of the dashed line, and $X_{(1)}$, represent irreducible sample confounding effects. Under i.i.d. sampling, this can be expressed as
@@
$=P(U_{(1)}\leq x)=F_{X_{(1)}},$
which uses the observation that $x$ is the minimum if and only if all $X$ are greater than $x$ [Gut, 2009].
@@
where $\Delta_{(k)}m$ and $\Delta_{(k)}q$ are the number of, respectively, observed and unobserved factors with positive probability in the sample’s $k$-th ordinal position or gap. These gap statistics are closely associated with the possible extensions among posets $\{G(X),G(Y)\}$ (such as those illustrated in Example. (1)).
Sample gap effect statistics will become relevant, in particular, because they reflect only the effect of new causes introduced within the gap ($\Delta y(u)$, $a<u_{k}<b$ for two sample factors $a$ and $b$ and unobserved causes $u_{k}\subseteq U$). Of particular interest are gap effects when in-sample factors are the same in subsequent effect observations. In this case, outcome differences in a gap $k$ reflect exclusively the effect of unobserved factors ’falling’ in $k$, $0<k\leq m$. In a square diagram with $X=\{a,b,c\}$, for example, $Y_{(k)}(X{=}a)-Y_{(k+1)}(X{=}a)$ denotes differences in outcome across its main diagonal, Fig.1(a). Eq.(5) suggest distinct levels of confounding across a sample’s effect observations,
@@
Effect observations of factor $a$ at $\textrm{argmin}_{k}X_{(k)}=\{a\}$ are a sample’s minimally confounded and maximally precise individual observations of $a$’s effect, $\textrm{Var}^{-1}[\Delta y(a)]$. In a single-row square, the effect of its first factor is observed at maximum precision. All other effects can be estimated confounded only by $\Delta_{(k)}y$, by conditioning on effect observations prior in the square row order (and their interaction effects).
@@
This suggests that, even when a factor $b$ is not observed at a minimal order $k=1$, unobserved confounding effects at $k=1$ can still be discounted from estimates by using other sample factor effects prior to $b$ in row order. For example, when $a$ is observed at $k=1$ but $b$ only at $k=2$, $\textrm{Var}[\Delta y_{(2)}(b\,|\,a)-\Delta y_{(1)}(a)]\leq\textrm{Var}[\Delta y_{(2)}(b\,|\,a)]$.
We will consider, in particular, the relationship between the number of permutations that $U{-}X$ systems can generate and their impact in the statistics of effects across sample gaps. According to Principle (1) the first is a requirement for the separability of causes and confounders, when using effect observations. Some distributions of unobserved causes across gaps can asymptotically generate all distinct orders (all permutations of unobserved factors) and some can’t. This can be described by a multinomial coefficient, or, by dividing the set of possible permutations in a $U{-}X$ system, $Pi_{n}(Z/\{a\})$, and the ideal full set of permutations $q!$,
@@
Figure 2: (a) sampling sequences in a sample, $X=\{a,b,c\}$, with confounder partition $U=\{\{u_{1}\},\{u_{2},u3\},\{u_{4}\}\}$, (b) normalized variation separation distance (distance to a Uniform distribution) of all backgrounds $\pi_{0}$ to other backgrounds after time $t$, vertical dashed lines mark limits on successive background randomness based on their inversions, Eq.(11) (top panel shows the same without a test of stationarity on gap distributions $p$), (c) randomization in graphical models with in-sample treatment $A$, other factors $W$, unobserved factors $U$, and outcome $Y$, (d) combinatorial relations in a square ($m=2$), circles are differences and gray circles are inter-unit factor value intersections (letters) between $x_{0}$ and other units.

### v6

@@ Abstract
The widely used ’Counterfactual’ definition of Causal Effects was derived for unbiasedness and accuracy - and not generalizability. We propose a Combinatorial definition for the External Validity (EV) of intervention effects. We first define the concept of an effect observation ’background’. We then formulate conditions for effect generalization based on samples’ sets of (observed and unobserved) backgrounds. This reveals two limits for effect generalization: (1) when effects of a variable are observed under all their enumerable backgrounds, or, (2) when backgrounds have become sufficiently randomized. We use the resulting combinatorial framework to re-examine several issues in the original counterfactual formulation: out-of-sample validity, concurrent estimation of multiple effects, bias-variance tradeoffs, statistical power, and connections to current predictive and explaining techniques.
Methodologically, the definitions also allow us to replace the parametric estimation problems that followed the counterfactual definition by combinatorial enumeration and randomization problems in non-experimental samples. We use the resulting non-parametric framework to demonstrate (External Validity, Unconfoundness and Precision) tradeoffs in the performance of popular supervised, explaining, and causal-effect estimators. We also illustrate how the approach allows for the use of supervised and explaining methods in non-i.i.d. samples. The COVID19 pandemic highlighted the need for learning solutions to provide predictions in severally incomplete samples. We demonstrate applications in this pressing problem.
@@ 1 Introduction
Non-parametric effect estimators (such as permutation tests  [Lee, 1990] and feature attribution methods  [Burkart and Huber, 2021]) can often only be used under strict assumptions of no unobserved factors. Consider, for example, a sample that contains units that both have received a treatment (’treatment’ units) and others that did not (’control’). To estimate effects, a simple strategy is to permute treatment assignment among treatment-control pairs that are similar enough across relevant variables  [Morgan and Winship, 2007, Rubin, 2005]. This strategy relies, however, on the assumption that paired units are exchangeable besides their treatment status. The assumption is easily violated if units differ in any unobserved way (even in high-dimensional samples it is hard for researchers to make the claim that there are no overlooked, relevant distinctions). The main challenge is that of unobserved confounding, and of sampling conditions where uncontrolled variation of out-of-sample factors bias in-sample estimates. Under these conditions, sample observation of effects, for example, reflect not only the true effect of sample factors, but also of uncontrolled and unobserved factors. In Experimental Design, randomization plays a crucial role in alleviating the effects of unobserved confounding  [Zhang and Zhao, 2023, Kempthorne and Doerfler, 1969]. Randomized Experiments provide a solid foundation, and are still the golden rule, for making causal inferences about the effects of treatments in the Sciences. Non-experimental samples can be seen as experimental samples where randomization did not occur perfectly. Despite these known shortcomings, permutation and Shapley-value based feature attribution methods  [Burkart and Huber, 2021, Ribeiro, 2022a] continue to be used in practice in non-experimental and partially observed data, with little clarity over the resulting effect estimates’ generalizability and biases. We provide a combinatorial definition for effect generalizability and demonstrate how to select subsamples in non-experimental data that lead to effect estimates with high generalizability and low confoundness. We demonstrate uses of these concepts when estimating causal effects non-parametrically or when explaining the performance of supervised predictors. The latter is currently an intensely studied subject in Machine Learning and Artificial Intelligence  [Lundberg and Lee, 2017, Burkart and Huber, 2021].
@@
A sample’s Data Generating Process (DGP) refers to the underlying mechanism or model that generated the sample. Partition the set $Z$ of DGP variables into two subsets, $Z=X_{m}\cup U_{q}$, the first with its $m$ observed variables $X_{m}=\{-1,+1\}^{m}$, and the second with $q$ unobserved $U_{q}=\{-1,+1\}^{q}$. Imagine also variables in $X$ and $U$ are constrained by precedence relationships, as described by two partially ordered set (posets) $G(Z)=\{G(X),G(U)\}$. The distinct ways in which $\{G(X),G(U)\}$ can be combined is described by the distinct linear extensions among these posets. For poset chains $\{G(X),G(U)\}=\{\;\mathbf{a}<\mathbf{b},\;\;u_{1}<u_{2}<u_{3}\;\}$, for example, the extensions are
@@
The different extensions can be seen as different confounding conditions under which sample factors $X$ can be observed. We imagine real-world samples can be collected under any such conditions, and we would like to formulate statistical properties for effects across partially observed samples, $q>0$. The likelihood of a particular $U{-}X$ extension follow, in turn, relationships between the DGP probability distributions for $X$ and $U$,
@@ 1.0.1 Alternative Orders and Confounding: A Preliminary illustration
For illustration, consider the effects observed in a sample with a true cause $a$, a confounder $b$ (with effects on $y$ fully correlated with its root cause $a$), and a spurious variable $c$, $X=\{a,b,c\}$. We will use Latin square diagrams (’squares’) as aid to illustrate several concepts in this article. Fig.1(a) shows a $4$-factor square. Each square is associated with a sample unit $x_{0}$, its reference unit, which stratifies the sample (i.e., other units) across its cells. A square with reference $x_{0}$ correspond to the set of effect observations with singleton factor differences,
$\begin{split}\bigcup_{x\in\mathcal{P}(X)}{{\Delta}}y(x-x_{0}=\{a\}),\qquad\forall\,a\in X.\end{split}$
where $\mathcal{P}(X)$ is the powerset (set of sets) of $X$ and ${\Delta}y(x-x_{0})$ is an observation of effect in an outcome-of-interest, $y\in\mathbb{R}$, after an intervention on factors $x-x_{0}$. A complete square contains therefore $\binom{m}{2}=\frac{m\cdot(m-1)}{2}$ effect observations, and a sample (or all its squares) is limited to $2^{m}$ observations. We will say that squares’ diagonals enumerate, for example, observations of effect for fixed factors across different (observed) ’backgrounds’. For the cause $a$ and confounder $b$, differences in effect observations across diagonals are
Example 3 (Iterative higher-order effects).
@@
We see that when the cause $a$ is in the sample, it sustains its effects throughout arbitrary sample imputations (indicated by the conditional statements), while the confounder also assumes negative values. In a square with factors $X$, observed under different orders, effect observations for confounders change many times, while for causes they remain invariant throughout the observed range of background variation.
@@
For any factors $a,b\in Z$, observing their effects under conditions $\Delta y(a\,|\,b)$ and $\Delta y(b\,|\,a)$, or under ordinal relations $a>b$ and $b>a$, is necessary to identify, from effect observations, whether $a$ confounds $b$ (and vice versa). When observed under such conditions, effect variances, $\textrm{Var}[\Delta\hat{y}(b)]$, for causes, confounders and spurious factors become distinct, and have expected values of 00, $\nicefrac{{\rho_{ab}}}{{4}}\Delta y(a)$, $\nicefrac{{1}}{{2}}\Delta y(a)$, where $\rho_{ab}$ is the Spearman’s rank correlation coefficient between confounder and its root cause, and $\Delta y(a)$ the root cause’s effect.
According to the principle, under specific ordinal conditions on $U{-}X$ factors, effect observations for confounders have an expected effect variance of $0.25\times\rho_{ab}$ (i.e., the number of times the confounder appears before its root cause in square rows) of the true effect $\Delta y(a)$ and causal variables 00. One such condition is of a completely observed sample, $q=0$. In this case, the principle is associated with the observation of each effect under the cyclic permutation of all others factors (as illustrated by distinct square rows). The principle states then that effect variances, ${\rm I\kern-1.99997ptE}[\,\textrm{Var}[\Delta\hat{y}(a)]\,]$, are a sufficient statistic to identify confounders under a fixed level of cause-confounder correlation $\rho_{ab}$. Conditions when $q>0$ are established below.
@@ 1.1 Effect estimates and Unobserved Sample Variation
Before formulating necessary conditions for the generalizability and confoundness of effect observation in generic samples, we formulate our working definition of effects, their generalizability, and a roadmap to the article. According to the Potential Outcomes framework definitions for what is a causal effect  [Morgan and Winship, 2007, Rubin, 2005], if $y$ is an outcome-of-interest and $a$ is a treatment indicator, then the causal effect of $a$ is the difference
@@
where $y_{\leq}$ is the observed outcome $y$ of a sample unit with, exclusively, the set of factors before $a$ in the permutation order $\pi$, and $y_{>}$ of units also containing the factors after $a$. Difference between units in Eq.(3) constitutes observations of effects of the same factor in distinct backgrounds, $\pi\in\Pi_{n}(Z/\{a\})$. Each background fixes a set of factors for their ’treatment’ (before $a$ in $\pi$, or, sample effects that do not reflect the effect of $a$) and ’control’ (after $a$) for sample units.
Furthermore, $\Pi_{n}(Z/\{a\})$ denotes a subset of the $(m+q)!$ full set of permutations of factors $Z$ possible between sequential observations of $a$. Which subsets can offer guarantees over the generalizability of sample effect estimates? We formulate the concept of External Validity increasing (EV-increasing) permutations as those that can increase the generalizability of effects in samples. The article’s main argument is structured as follows,
@@
EV is here the inverse of the variance (i.e., the precision) of observed effects under all (e.g., asymptotically) EV-increasing backgrounds, $\Pi_{n}(Z/\{a\})$, for a treatment $a$. This defines effects in a way that is almost opposite to Eq. (1). It calls for effects to be observed under large variation, as opposed to no variation. According to Eq.(3,4), the counterfactual ideal in Eq.(1) is, in fact, the estimate with minimal EV. Underlying these definitions and Principle (1) is the observation that, unlike confounders, causes can be fundamentally defined as ’perfect controls’  [Pearl, 2000]. Similar definitions have been explored in model robustness research  [Peters et al., 2016, Buehlmann, 2020]. The non-parametric definitions here frame the problem, in particular, not as simply the invariance (of effects), but invariance of effects under maximum variance of their ’backgrounds’. Shifting the focus to the later, $\Pi_{n}(Z/\{a\})$, is a key but challenging part of the proposed solution.
Figure 1: (a) $4{\times}4$ Latin-Square (’square’) as sets of effect observations for a sample unit $x_{0}\subseteq X$, (b) $3$ samples (horizontal lines, top) with increasing dimension $m=\{1,2,4\}$ and their ordinal statistics (interval ticks) , as well as alternative samples of confounding unobserved factors $U$ (circles) (top), resulting distribution and percentage of enumerable permutations of effect observations across ordinal variables (bottom and right).
@@ 2 Ordinal Statistics and Sample Confounding Gaps
If $\{X_{a},X_{b},\ldots,X_{m}\}$ are random samples from a particular DGP distribution of $X_{m}$, it will be important to know whether occurrences of a factor $a$ are expected to follow or precede occurrences of another factor $b$. The likelihood of such order relations is determined by the DGP, as illustrated in Example (2). To describe such patterns more formally, we define $X_{(1)},X_{(2)},\ldots,X_{(m)}$ to be an ordered version of sample factors  [Gut, 2009], that is,
@@
Notice that $X_{k}$ are random variables, $0<k\leq m$. For simplicity, we assume that, in sufficiently large samples and time resolution, ranks do not coincide, leading to strict inequalities (as opposed to $\leq$).
Suppose that $X_{1},X_{2},\ldots,X_{m}$ follow a continuous distribution with PDF $P$. We will use $X$, $U$ or $Z$ as alternative labels when considering distributions over observed, unobserved or either factors. Imagine then a random sample of $U$ factors whose minimum is below an in-sample value $x$. Fig.1(b) give examples where circles are $U$ samples and interval ticks are $X$ samples. The successive lines show $X_{(k)}$ in samples with increasing dimensionality $m$. Circles to the left of the dashed line, and $X_{(1)}$, introduce irreducible unobserved confounding effects to the sample. Under i.i.d. sampling, this can be expressed as
@@
$=P(U_{(1)}\leq x)$
@@
where $\Delta_{(k)}m$ and $\Delta_{(k)}q$ are the number of, respectively, observed and unobserved factors with positive probability in the sample’s $k$-th ordinal position or gap. These gap statistics are closely related to the possible extensions among posets $\{G(X),G(Y)\}$ (such as those illustrated in Example (1)).
Sample gap effect statistics will become relevant because they reflect only the effect of new causes introduced within the gap ($\Delta y(u)$, $a<u_{k}<b$ for two sample factors $a$ and $b$ and unobserved causes $u_{k}\subseteq U$). Of particular interest are gap effects when in-sample factors are the same in subsequent effect observations. In this case, outcome differences in a gap $k$ reflect exclusively the effect of unobserved factors ’falling’ in $k$, $0<k\leq m$. In a square diagram with $X=\{a,b,c\}$, for example, $Y_{(k)}(X{=}a)-Y_{(k+1)}(X{=}a)$ denotes differences in outcome across its main diagonal, Fig.1(a). Eq.(5) thus suggests distinct levels of confounding across a sample’s effect observations,
@@
Effect observations of factor $a$ at $\textrm{argmin}_{k}X_{(k)}=a$ are a sample’s minimally confounded and maximally precise individual observations of $a$’s effect, $\textrm{Var}^{-1}[\Delta y(a)]$. In a single-row square, the effect of its first factor is observed at maximum precision. All other effects can be estimated confounded only by $\Delta_{(k)}y$, by conditioning on effect observations prior in the square row order (and their interaction effects).
@@
The first proposition suggests that, even when a factor $b$ is not observed at a minimal order $k=1$, unobserved confounding effects at $k=1$ can still be discounted from estimates by using other sample factor effects prior to $b$ in row order. For example, when $a$ is observed at $k=1$ but $b$ only at $k=2$, $\textrm{Var}[\Delta y_{(2)}(b\,|\,a)-\Delta y_{(1)}(a)]\leq\textrm{Var}[\Delta y_{(2)}(b\,|\,a)]$.
We will consider, in particular, the relationship between the number of permutations that $U{-}X$ systems can generate and their impact in the statistics of effects across sample gaps. According to Principle (1) the first is a requirement for the separability of causes and confounders, when using effect observations. Some distributions of unobserved causes across gaps can asymptotically generate all distinct orders (all permutations of unobserved factors) and some can’t. This can be described by a multinomial coefficient, or, by dividing the set of possible permutations in a $U{-}X$ system, $\Pi_{n}(Z/\{a\})$, and the ideal full set of $q!$ permutations of $U$,
@@
Figure 2: (a) sampling sequences in a sample, $X=\{a,b,c\}$, with confounder partition $U=\{\{u_{1}\},\{u_{2},u3\},\{u_{4}\}\}$, (b) normalized variation separation distance (distance to a Uniform distribution) of all backgrounds $\pi_{0}$ to other backgrounds after time $t$, vertical dashed lines mark limits on successive background randomness based on their number of inversions, Eq.(11) (top panel shows the same without a test of stationarity on gap distributions $p$), (c) randomization in graphical models with in-sample treatment $A$, other factors $W$, unobserved factors $U$, and outcome $Y$, (d) combinatorial relations in a square ($m=2$), circles are differences and gray circles are inter-unit factor value intersections (letters) between $x_{0}$ and other units.